Obtain a solution set for the linear system

x1-2x2-3x3=0

-2x1+4x2+6x3=0

X1+2x2-5=0


1
Expert's answer
2021-02-04T01:39:57-0500

Obtain a solution set for the linear system

x1 - 2x2 - 3x3=0

-2x1+4x2+6x3=0

X1+2x2-5x3=0


"\begin{pmatrix}
 1 & -2 &-3\\-2 &4 &6\\1 &2 &-5
\end{pmatrix}" "\begin{pmatrix}
 x_1 \\
 x_2 \\ x_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\0
\end{pmatrix}"


"\begin{pmatrix}
 1 & -2 &-3&| &0\\-2 &4 &6&|&0\\1 &2 &-5&|&0
\end{pmatrix}R_3
 \mapsto2R_3 +R_2" "\begin{pmatrix}
 1 & -2 &-3&| &0\\-2 &4 &6&|&0\\0 &8 &-4&|&0
\end{pmatrix}"



"\begin{pmatrix}
 1 & -2 &-3&| &0\\-2 &4 &6&|&0\\0 &8 &-4&|&0
\end{pmatrix} R_2 \mapsto R_2 +2R_1" "\begin{pmatrix}
 1 & -2 &-3&| &0\\0 &0 &0&|&0\\0 &8 &-4&|&0
\end{pmatrix}"


"x_1-2x_2-3x_3=0........(i)"


"8x_2-4x_3=0..........................(ii)"

from equation (ii)

"8x_2=4x_3 \implies x_3=2x_2"

we can view as determining x3 in term of x2 , with no restriction placed on x2 a all

"x_2" ="\alpha" (arbitrary i.e. "\alpha" can be any number)

then "x_3=2\alpha"


put x2 and x3 into eqn(i)

"x_1-2(\alpha)-3(2\alpha)=0"


"x_1=8\alpha"

"\because" the linear system contains one free parameter